# Ask, and it shall be given: On the Turing completeness of prompting

### v2

@@ 1 Introduction
Nevertheless, existing theoretical studies fail to align with the LLM prompting practice (i.e., the one-model-many-tasks paradigm). In the era of LLMs, people train a single general-purpose LLM and provide the LLM with different prompts to perform different tasks. Since the success of GPT (Brown et al., 2020), the LLM prompting paradigm has revolutionized machine learning (Liu et al., 2023b). For example, a bonus capability arising from prompting is zero-shot learning (Wei et al., 2022a): when provided with suitable prompts, LLMs can even perform novel tasks that are not present in their training corpora. Such empirical success calls for a theoretical understanding of the LLM prompting paradigm:
@@
  • Simple construction. In fact, it is not hard to deduce the existence of such a Transformer $\varGamma$ by combining Theorem 1 of Hennie & Stearns (1966) and Theorem 3.4 of Pérez et al. (2019), but explicitly constructing the Transformer via that approach is cumbersome. Instead, this paper provides a simple constructive proof, which makes it easy to further study the properties of the construction such as CoT complexity and precision complexity.
@@ 1.2 Technical overview
Our constructed $\varGamma$ can compute any $\mathsf{TIME}(t(n))$ function within $\mathrm{O}(t(n)\log t(n))$ CoT steps.
@@
Our constructed $\varGamma$ can compute any $\mathsf{TIME}(t(n))$ function within $\mathrm{O}(\log(n+t(n)))$ bits of precision.
@@ 2 Preliminaries
Let $ϵ\upepsilon$ denote the empty string. Given a finite alphabet $\varSigma$, let $\varSigma^{n}$ ($n\geq 0$) denote the set of length-$n$ strings over $\varSigma$, let $\varSigma^{*}:=\bigcup_{n\geq 0}\varSigma^{n}$ denote the set of all finite strings over $\varSigma$, let $\varSigma^{+}:=\bigcup_{n\geq 1}\varSigma^{n}$ denote the set of all non-empty finite strings over $\varSigma$, and let $\varSigma^{\upomega}$ denote the set of all countably infinite strings over $\varSigma$. Each symbol in $\varSigma$ is called a token. For a string $\boldsymbol{x}\in\varSigma^{*}$, let $|\boldsymbol{x}|$ denote the length of the string, and let $\boldsymbol{x}^{n}$ denote repeating the string $n$ times. For two strings $\boldsymbol{x},\boldsymbol{y}\in\varSigma^{*}$, let $\boldsymbol{x}\cdot\boldsymbol{y}$ denote their concatenation. Given a string $\boldsymbol{x}=x_{0}\cdot x_{1}\cdots x_{|\boldsymbol{x}|-1}\in\varSigma^{*}$, for an index $i\in\mathbb{Z}$, let $x_{i}$ denote the $(i\bmod|\boldsymbol{x}|)$-th token of $\boldsymbol{x}$; for indices $i,j\in\mathbb{Z}$ with $i\bmod|\boldsymbol{x}|\leq j\bmod|\boldsymbol{x}|$, let $\boldsymbol{x}_{i:j}$ denote the substring $x_{i\bmod|\boldsymbol{x}|}\cdot x_{(i\bmod|\boldsymbol{x}|)+1}\cdots x_{(j\bmod|\boldsymbol{x}|)-1}$.
@@ 2.2 Neural networks
Decoder-only Transformers. Mainstream LLMs are based on decoder-only Transformers (Radford et al., 2018). Given a finite token alphabet $\varSigma$, a decoder-only Transformer $\varGamma:\varSigma^{+}\to\varSigma$ (with greedy decoding) is a composition $\varGamma:=\operatorname*{\arg\max}\circ\varGamma_{\text{out}}\circ\varGamma_{L-1}\circ\cdots\circ\varGamma_{0}\circ\varGamma_{\text{emb}}$ consists of an embedding layer $\varGamma_{\text{emb}}:\varSigma\to\mathbb{R}^{d}$, causal-attention Transformer layer $\varGamma_{l}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ ($l=0,\dots,L-1$), and an output layer $\varGamma_{\text{out}}:\mathbb{R}^{d}\to\mathbb{R}^{\varSigma}$. A token embedding is a map $\operatorname{emb}:\varSigma\to\mathbb{R}^{d}$, and a positional encoding is a map $\operatorname{pos}:\mathbb{N}\to\mathbb{R}^{d}$. Following the common practice (Vaswani et al., 2017), given the input token sequence $\boldsymbol{v}=v_{0}\cdots v_{|\boldsymbol{v}|-1}\in\varSigma^{+}$, the embedding layer $\varGamma_{\text{emb}}$ adds a positional encoding $\operatorname{pos}(i)$ to each token embedding $\operatorname{emb}(v_{i})$:
@@
where $\operatorname{pos}$ is a computable function. Following existing theoretical works (e.g., Pérez et al., 2019; Hao et al., 2022; Merrill & Sabharwal, 2024a), we use $\operatorname{hardmax}$ attention as a realistic abstraction of $\operatorname{softmax}$. Given an $\mathbb{R}$ sequence $\boldsymbol{s}$, if the maximum value of $\boldsymbol{s}$ appears $t$ times, then we define $\operatorname{hardmax}_{i}(\boldsymbol{s}):=\frac{1}{t}$ if $s_{i}$ equals the maximum value and $\operatorname{hardmax}_{i}(\boldsymbol{s}):=0$ otherwise. Each Transformer layer $\varGamma_{l}$ has $H_{l}$ attention heads followed by a ReLU neural network $\phi_{l}:\mathbb{R}^{d}\to\mathbb{R}^{d}$. Each attention head $k$ has a query map $\operatorname{qry}_{l,k}:\mathbb{R}^{d}\to\mathbb{R}^{d_{l,k}}$, a key map $\operatorname{key}_{l,k}:\mathbb{R}^{d}\to\mathbb{R}^{d_{l,k}}$, a value map $\operatorname{val}_{l,k}:\mathbb{R}^{d}\to\mathbb{R}^{d}$, and a similarity map $\operatorname{sim}_{l,k}:\mathbb{R}\to\mathbb{R}$ ($k=0,\dots,H_{l}-1$), all of which are ReLU neural networks. The similarity scores between two tokens $v_{i}$ and $v_{j}$ are
@@
Given a Transformer $\varGamma$, let $\operatorname{generate}_{\varGamma}:\varSigma^{+}\to\varSigma^{+}\cup\varSigma^{\upomega}$ denote autoregressive generation using $\varGamma$. Specifically, let $\boldsymbol{w}\in\varSigma^{+}$ denote the input sequence and $\boldsymbol{y}\in\varSigma^{*}$ the (current) output sequence. Initially, $\boldsymbol{y}$ is an empty string. In each generation step, we append $\varGamma(\boldsymbol{v}\cdot\boldsymbol{y})$ to $\boldsymbol{y}$. The generation process ends once the last token of $\boldsymbol{y}$ is a so-called stop token \$$\texttt{\$}\in\varSigma$, and we define $\operatorname{generate}_{\varGamma}(\boldsymbol{w}):=\boldsymbol{y}$ at last. It is possible that generation never ends, in which case we have $\boldsymbol{y}\in\varSigma^{\upomega}$. The autoregressive generation procedure is formally presented in Algorithm 1. A Transformer is said to use log-precision for a computable function $\varphi$ if the intermediate computation of all generation steps uses $\mathrm{O}(\log n)$ bits of precision for every length-$n$ input of function $\varphi$ (Merrill & Sabharwal, 2023).
@@ 3 Turing completeness of prompting
Here, $\varSigma,\varGamma,\operatorname{tokenize},\operatorname{readout}$ are independent of $\varphi$; $\boldsymbol{\pi}_{\varphi}$ is independent of $\boldsymbol{x}$; for any input $\boldsymbol{x}\in\{\mathtt{0},\mathtt{1}\}^{*}$, $\operatorname{tokenize}$ and $\operatorname{readout}$ run in $\mathrm{O}(|\boldsymbol{x}|)$ and $\mathrm{O}(|\varphi(\boldsymbol{x})|)$ time on a random access machine (RAM), respectively.
@@ 3.1 Construction of prompts
$P(j,\iota_{j}):=\begin{cases}\tau{\texttt{!}}\cdot\texttt{-}^{j-k}\cdot\texttt{@}&\text{if }\iota_{j}=\tau{\texttt{!}}_{k}\text{ and }k<j;\\ \tau{\texttt{!}}\cdot\texttt{+}^{k-j}\cdot\texttt{@}&\text{if }\iota_{j}=\tau{\texttt{!}}_{k}\text{ and }k>j;\\ \tau{\texttt{?}}\cdot\texttt{-}^{j-k}\cdot\texttt{@}&\text{if }\iota_{j}=\tau{\texttt{?}}_{k}\text{ and }k<j;\\ \tau{\texttt{?}}\cdot\texttt{+}^{k-j}\cdot\texttt{@}&\text{if }\iota_{j}=\tau{\texttt{?}}_{k}\text{ and }k>j;\end{cases}$ (9)
@@ 3.4 Construction of Transformer
Therefore, we can retrieve the smallest $i$ with $\boldsymbol{u}_{i}^{\mathsf{T}}\boldsymbol{u}_{|\boldsymbol{v}|-1}=1$ using query vector $(\boldsymbol{u}_{|\boldsymbol{v}|-1},p_{|\boldsymbol{v}|-1})^{\mathsf{T}}$ and key vector $(\boldsymbol{u}_{i},\frac{1}{i+1})^{\mathsf{T}}$ in causal attention.
@@ 4.1 Efficient simulation of TMs by $2$-PTMs
Theorem 4.1 shows that our $2$-PTMs are Turing-complete and nearly as efficient as TMs. To prove it, we need the following Lemma 4.2 to establish an equivalence between two-tape TMs and $2$-PTMs.
@@
Proof sketch of Lemma 4.2.
For any computable function $\varphi\in\mathsf{TIME}_{2}(t(n))$, there exists a two-tape TM $M=(Q,q_{\text{start}},q_{\text{halt}},\{\mathtt{0},\mathtt{1}\},\textvisiblespace=\mathtt{0},m=2,\delta)$ that computes $S(\varphi(\boldsymbol{x}))$ from $S(\boldsymbol{x})$ within time $\mathrm{O}(t(n))$, according to Shannon (1956). Suppose w.l.o.g. that $Q=\{0,1,\dots,K\}$ ($K\in\mathbb{N}_{+}$) and that $q_{\text{start}}=0$, $q_{\text{halt}}=K$. We will construct a $2$-PTM $\boldsymbol{\iota}$ of length $|\boldsymbol{\iota}|=27K+1$ that simulates $M$ within time $\mathrm{O}(t(n))$.

### v3

@@ 1 Introduction
Nevertheless, existing theoretical studies fail to align with the LLM prompting practice (i.e., one-model-many-tasks). In the era of LLMs, people train a single general-purpose LLM and provide the LLM with different prompts to perform different tasks. Since the success of GPT (Brown et al., 2020), the LLM prompting paradigm has revolutionized machine learning (Liu et al., 2023b). For example, a bonus capability arising from prompting is zero-shot learning (Wei et al., 2022a): when provided with suitable prompts, LLMs can even perform novel tasks not present in their training corpora. Such empirical success calls for a theoretical understanding of the LLM prompting paradigm:
@@
  • Simple construction. In fact, it is not hard to deduce the existence of such a Transformer $\varGamma$ by combining Theorem 1 of Hennie & Stearns (1966) and Theorem 3.4 of Pérez et al. (2019), but explicitly constructing the Transformer via that approach is cumbersome. Instead, we provide a simple constructive proof, which makes it easy to further study the properties of the construction such as CoT complexity and precision complexity.
@@ 1.2 Technical overview
Our constructed $\varGamma$ can compute any $\mathsf{TIME}(t(n))$ function within $\mathrm{O}(t(n)\log t(n))$ CoT steps. ∎
@@
Our constructed $\varGamma$ can compute any $\mathsf{TIME}(t(n))$ function within $\mathrm{O}(\log(n+t(n)))$ bits of precision. ∎
@@ 2 Preliminaries
Let $ϵ\upepsilon$ denote the empty string. Given an alphabet $\varSigma$, let $\varSigma^{n}$ ($n\geq 0$) denote the set of length-$n$ strings over $\varSigma$, let $\varSigma^{*}:=\bigcup_{n\geq 0}\varSigma^{n}$ denote the set of all finite strings over $\varSigma$, let $\varSigma^{+}:=\bigcup_{n\geq 1}\varSigma^{n}$ denote the set of all non-empty finite strings over $\varSigma$, and let $\varSigma^{\upomega}$ denote the set of all countably infinite strings over $\varSigma$. Each symbol in $\varSigma$ is called a token. For a string $\boldsymbol{x}\in\varSigma^{*}$, let $|\boldsymbol{x}|$ denote the length of the string, and let $\boldsymbol{x}^{n}$ denote repeating the string $n$ times. For two strings $\boldsymbol{x},\boldsymbol{y}\in\varSigma^{*}$, let $\boldsymbol{x}\cdot\boldsymbol{y}$ denote their concatenation. Given a string $\boldsymbol{x}=x_{0}\cdot x_{1}\cdots x_{|\boldsymbol{x}|-1}\in\varSigma^{*}$, for an index $i\in\mathbb{Z}$, let $x_{i}$ denote the $(i\bmod|\boldsymbol{x}|)$-th token of $\boldsymbol{x}$; for indices $i,j\in\mathbb{Z}$ with $i\bmod|\boldsymbol{x}|\leq j\bmod|\boldsymbol{x}|$, let $\boldsymbol{x}_{i:j}$ denote the substring $x_{i\bmod|\boldsymbol{x}|}\cdot x_{(i\bmod|\boldsymbol{x}|)+1}\cdots x_{(j\bmod|\boldsymbol{x}|)-1}$.
@@ 2.2 Neural networks
Decoder-only Transformers. Mainstream LLMs are based on decoder-only Transformers (Radford et al., 2018). Given a finite token alphabet $\varSigma$, a decoder-only Transformer $\varGamma:\varSigma^{+}\to\varSigma$ (with greedy decoding) is a composition $\varGamma:=\operatorname*{\arg\max}\circ\varGamma_{\text{out}}\circ\varGamma_{L-1}\circ\cdots\circ\varGamma_{0}\circ\varGamma_{\text{emb}}$ consists of an embedding layer $\varGamma_{\text{emb}}:\varSigma^{+}\to(\mathbb{R}^{d})^{+}$, causal-attention Transformer layer $\varGamma_{l}:(\mathbb{R}^{d})^{+}\to(\mathbb{R}^{d})^{+}$ ($l=0,\dots,L-1$), and an output layer $\varGamma_{\text{out}}:\mathbb{R}^{d}\to\mathbb{R}^{\varSigma}$. A token embedding is a map $\operatorname{emb}:\varSigma\to\mathbb{R}^{d}$, and a positional encoding is a map $\operatorname{pos}:\mathbb{N}\to\mathbb{R}^{d}$. Following the common practice (Vaswani et al., 2017), given the input token sequence $\boldsymbol{v}=v_{0}\cdots v_{|\boldsymbol{v}|-1}\in\varSigma^{+}$, the embedding layer $\varGamma_{\text{emb}}$ adds a positional encoding $\operatorname{pos}(i)$ to each token embedding $\operatorname{emb}(v_{i})$:
@@
where $\operatorname{pos}$ is a computable function. Following existing theoretical works (e.g., Pérez et al., 2019; Hao et al., 2022; Merrill & Sabharwal, 2024a), we use $\operatorname{hardmax}$ attention as a realistic abstraction of $\operatorname{softmax}$. Given an $\mathbb{R}$ sequence $\boldsymbol{s}$, if the maximum value of $\boldsymbol{s}$ appears $t$ times, then one has $\operatorname{hardmax}_{i}(\boldsymbol{s}):=\frac{1}{t}$ if $s_{i}$ equals the maximum value and $\operatorname{hardmax}_{i}(\boldsymbol{s}):=0$ otherwise. Each Transformer layer $\varGamma_{l}$ has $H_{l}$ attention heads followed by a ReLU neural network $\phi_{l}:\mathbb{R}^{d}\to\mathbb{R}^{d}$. Each attention head $k$ has a query map $\operatorname{qry}_{l,k}:\mathbb{R}^{d}\to\mathbb{R}^{d_{l,k}}$, a key map $\operatorname{key}_{l,k}:\mathbb{R}^{d}\to\mathbb{R}^{d_{l,k}}$, a value map $\operatorname{val}_{l,k}:\mathbb{R}^{d}\to\mathbb{R}^{d}$, and a similarity map $\operatorname{sim}_{l,k}:\mathbb{R}\to\mathbb{R}$ ($k=0,\dots,H_{l}-1$), all of which are ReLU neural networks. The similarity scores between two tokens $v_{i}$ and $v_{j}$ are
@@
Given a Transformer $\varGamma$, let $\operatorname{generate}_{\varGamma}:\varSigma^{+}\to\varSigma^{+}\cup\varSigma^{\upomega}$ denote autoregressive generation using $\varGamma$. Specifically, let $\boldsymbol{w}\in\varSigma^{+}$ denote the input sequence and $\boldsymbol{y}\in\varSigma^{*}$ the (current) output sequence. Initially, $\boldsymbol{y}$ is an empty string. In each generation step, we append $\varGamma(\boldsymbol{w}\cdot\boldsymbol{y})$ to $\boldsymbol{y}$. The generation process ends once the last token of $\boldsymbol{y}$ is a so-called stop token \$$\texttt{\$}\in\varSigma$, and we define $\operatorname{generate}_{\varGamma}(\boldsymbol{w}):=\boldsymbol{y}$ at last. It is possible that generation never ends, in which case we have $\boldsymbol{y}\in\varSigma^{\upomega}$. The autoregressive generation procedure is formally presented in Algorithm 1. A Transformer is said to use log-precision for a computable function $\varphi$ if the intermediate computation of all generation steps uses $\mathrm{O}(\log n)$ bits of precision for every length-$n$ input of function $\varphi$ (Merrill & Sabharwal, 2023).
@@ 3 Turing completeness of prompting
Here, $\varSigma,\varGamma,\operatorname{tokenize},\operatorname{readout}$ are independent of $\varphi$; $\boldsymbol{\pi}_{\varphi}$ is independent of $\boldsymbol{x}$; for any input $\boldsymbol{x}\in\{\mathtt{0},\mathtt{1}\}^{*}$, $\operatorname{tokenize}$ and $\operatorname{readout}$ run in $\mathrm{O}(|\boldsymbol{x}|)$ and $\mathrm{O}(|\varphi(\boldsymbol{x})|)$ time on a RAM, respectively.
@@ 3.1 Construction of prompts
$P(j,\iota_{j}):=\begin{cases}\sigma\cdot\texttt{-}^{j-k}\cdot\texttt{@}&\text{if }\iota_{j}=\sigma_{k}\text{ and }k<j,\\ \sigma\cdot\texttt{+}^{k-j}\cdot\texttt{@}&\text{if }\iota_{j}=\sigma_{k}\text{ and }k>j,\end{cases}\qquad\sigma\in\{\tau\textnormal{{!}},\tau\textnormal{{?}}\}_{\tau\in\{\mathtt{A},\mathtt{B}\}},$ (9)
@@ 3.4 Construction of Transformer
Therefore, we can retrieve the smallest $i$ with $\boldsymbol{u}_{i}^{\mathsf{T}}\boldsymbol{u}_{|\boldsymbol{v}|-1}=1$ by using query vector $(\boldsymbol{u}_{|\boldsymbol{v}|-1},p_{|\boldsymbol{v}|-1})^{\mathsf{T}}$ and key vector $(\boldsymbol{u}_{i},\frac{1}{i+1})^{\mathsf{T}}$ in causal attention.
@@ 4.1 Efficient simulation of TMs by $2$-PTMs
Theorem 4.1 shows that our $2$-PTMs are Turing-complete and nearly as efficient as TMs. To prove it, we need the following Lemma 4.2 to establish a relation between two-tape TMs and $2$-PTMs.
@@
Proof of Lemma 4.2.
For any computable function $\varphi\in\mathsf{TIME}_{2}(t(n))$, there is a two-tape TM $M=(Q,q_{\text{start}},q_{\text{halt}},\{\mathtt{0},\mathtt{1}\},\textvisiblespace=\mathtt{0},m=2,\delta)$ that maps $S(\boldsymbol{x})$ to $S(\varphi(\boldsymbol{x}))$ within time $\mathrm{O}(t(n))$, according to Shannon (1956). Suppose w.l.o.g. that $Q=\{0,1,\dots,K\}$ ($K\in\mathbb{N}_{+}$) and that $q_{\text{start}}=0$, $q_{\text{halt}}=K$. We will construct a $2$-PTM $\boldsymbol{\iota}$ of length $|\boldsymbol{\iota}|=27K+1$ that simulates $M$ within time $\mathrm{O}(t(n))$.
@@ 5 Conclusion
Acknowledgments
This work is supported by NSF (2416070). The content of the information in this document does not necessarily reflect the position or the policy of the Government, and no official endorsement should be inferred. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation here on.